After 5 matches the mean number of goals scored by a football team per match is 1.8. if they score 3 goals in their 6th match, what is the mean after the 6th Match
1
Expert's answer
2020-03-19T13:09:02-0400

After 5 games the team scored "1.8*5=9" goals.

After 6 games the team scored "9+3=12" goals.

The mean after 6 games is "\frac{12}{6}=2" goals.